A function whose output is always the same regardless of
the input is a constant function. Give several examples of
constant functions from real life.

Answer by Theo(13342)

the function f(x) = 0 * x will always yield an answer of 0.

the function of f(x) = x^0 will always yield an answer of 1.

the function of f(x) = 0 / anything will always yield an answer of 0.

the function of f(x) = anything / 0 will always yield an answer of infinity.

since infinity is undefined, the answer will always be undefined.

the function of f(x) = x/x will always yield an answer of 1.

these are the most common that i know of.